# Ricci-Notation Tensor Framework for Model-based Approaches to Imaging

### v3

@@ I Introduction
“Tensors” easily associate with learning-based approaches to imaging. After all, TensorFlow from Google Brain [1] and Tensor Comprehensions from Facebook AI [2] are machine learning systems connected with a well known deep learning revolution in object detection and classification. Nonetheless, imaging researchers like Mai et al. [3] highlight the importance of model-based tensor approaches, emphasizing explanatory value. Mai et al. also argue that a model-based approach helps to reduce the training data requirements for a learning-based approach, yielding what is in effect a hybrid approach.
@@
Imaging researchers are finding advantages of model-based tensor approaches in a variety of applications. These include: multi-exposure, multi-focus, and hyper/multi-spectral image fusion to increase dynamic range, depth-of-field, and spatial resolution [3, 4, 5]; inpainting, or tensor completion, of missing or corrupt pixels in images and videos, even rows or columns of the same [6, 7]; deblurring, or image restoration, posed as a constrained inverse problem often with added noise [8, 9, 10]; compressive sensing, a mixture of computational imaging and image compression, for light field imaging (LFI), synthetic aperature radar, and hyperspectral imaging modalities [11, 12, 13]; other constrained image enhancements, like ones that enforce 3D spherical invariances [14]; and image quality assessments, like ones tailored for LFI [15].
@@
Model-based imaging research involving tensors [3][15] exhibits patterns. Papers typically have a section or subsection, on fundamentals, that addresses unary, binary, and $N$-ary operations expressible with tensors, like multi-way or multi-index contractions, Kronecker and Khatri-Rao products, and Tucker, canonical polyadic (CP), and tensor trains (TT) constructs. More fundamental than operations, authors discuss an $n$-mode notation, especially for decomposition constructs, an index notation, also called Einstein notation, or a hybrid notation thereof. Sometimes, extended matrix-vector (EMV) operators are employed to represent Hadamard, or entrywise, and Kronecker, or outer, products. Notation can resemble MATLAB or Python+NumPy syntax for multidimensional arrays. Instead of matrix-vector (MV) or EMV algebra, authors prefer tensor algebra to represent and manipulate high-order image relationships, including generalizations of the singular value decomposition (SVD) for rank reduction and feature extraction, akin to principal components analysis, and optimizations where images are the unknown variables in a functional that suffers upon vectorization of variables.
@@
When literature on learning-based [1][2] and model-based [3][15] approaches to imaging use the word, “tensor” mainly or solely means a numeric data structure, usually an MDA, together with numeric operations, like those identified with overloadable operators in object-oriented programming languages. This paper uses the word likewise, avoiding requirements for deeper tensor character addressed, for example by Synge and Schild [16], in books on geometric tensor calculus.
@@
Whereas tensor notations and operations are in some senses implementation agnostic, a brief survey of tensor software is in order. Because its components, like the Cyclops Tensor Framework (CTF) [17, 18], may apply to tensor contractions, completions, and decompositions for imaging, the NorthWest Chemistry (NWChem) software ecosystem is noteworthy [19, 20]. With the Tensor Contraction Engine (TCE) and the Tensor Algebra for Many-Body Methods (TAMM) [21, 22], this ecosystem targets massively parallel computing and large-scale molecular simulations. Intended more for imaging and for desktop parallel computing, Tensor Learning in Python (TensorLy) offers tensor decompositions to improve efficiency, robustness, and explainability of deep learning networks [23, 24]. For desktop computing with MATLAB, the Partial Least Squares (PLS) and Multivariate Image Analysis (MIA) Toolboxes, cited in biomedical imaging research [25, 26], offer tensor decompositions and imaging customizations.
@@
As Harrison and Joseph argue [27], the codesign of tensor algebra and software, what they call a tensor framework, has benefits. Codesign facilitates programmatic and computational efficiencies for solving a variety of model-based problems. In support, they cite Bader and Kolda’s [28, 29] development of $n$-mode+ notation and the MATLAB Tensor Toolbox (MTT), also called the Tensor Toolbox for MATLAB. While Bader and Kolda contributed notation for decomposition constructs along with original software, others leveraged existing notations while contributing decomposition constructs and MATLAB toolboxes. These include structured data fusion (SDF) with Tensorlab, led by Sorber and De Lathauwer [30, 31], and the TT-Toolbox with the TT construct, by Osledets [32].
@@
This paper proposes the Ricci-notation tensor (RT) framework for model-based approaches to imaging. A successor to Harrison and Joseph’s numeric tensor (NT) framework [27], itself a complement to an EMV framework, the RT framework inherits their advantages. The proposed framework comprises RT algebra and RT software, defined as the RTToolbox [33], developed with this paper, plus MATLAB. The RT software parses the RT algebra, dispatching calculations to the MATLAB kernel, with no dependencies on other toolboxes.
@@
The NT software had options [27]. With LibNT, C/C++ end users expressed the NT algebra with programmatic and, especially, computational efficiency. With NTToolbox, comprising M-file and MEX wrappers for LibNT, MATLAB end users expressed the NT algebra with some programmatic and some computational penalties. The RT software has only the RTToolbox option [33], although this paper elaborates on its design in a C/C++-compatible way. Comprising just M-files, no compilation or linking is required. As per MATLAB Central recommendations [34], avoidance of MEX files and binaries means that sharing is easy via File Exchange.
For its computational efficiency, the RT software leverages compiled unary and binary pagewise functions introduced, in 2020 and 2022, to the MATLAB kernel. Of note, the kernel pagewise multiplication [35] resembles the LibNT lattice multiplication [27] introduced, in 2016, with the NT software. This paper elaborates on the lattice concept to facilitate RT left- and right-division using kernel functions. It also introduces pagewise concatenation, an interpreted function the RT software exploits to support $N$-ary tensor concatenation. Programmatic efficiency of the RT software is illustrated via software expressions cross-referenced to algebraic ones.
An image enhancement example, developed here, illustrates the RT algebra. Inspired by Sirbu et al.’s model-based research [36, 37, 38] for coronagraphy [39], the example adopts a simple coronagraph model with two planes related by 2D discrete Fourier transforms and a phase aberration. Starting with a Creative Commons public domain (CC0) image [40], called Airy_disk_D65, the impact of the model on a synthetic ground-truth image having an occulted star with nearby exoplanets is computed. The exoplanets become visible after correction of the aberration, treated as unknown.
@@ II-B Nonscalars and Broadcasting
Unlike the NT algebra, the RT algebra supports outer additions, subtractions, concatenations, and relations. The unifying idea is that any tensor of one degree may be expanded to an equivalent tensor of higher degree via an outer product with a unit tensor of the difference degree. Consider the following equations and relation, which consistently employ a degree-two matrix, $\mathbf{A}$, four degree-one column-vectors, $\mathbf{1}$, $\mathbf{b}$, $\mathbf{c}$, and $\mathbf{1}$, and an entrywise absolute-value operation, ||$\cdot$|| ⋅ |:
@@
$\alpha_{\ell}^{j}$ =$=\begin{cases}1,&\ell=j\text{,}\\ 0,&\text{otherwise}\text{,}\end{cases}$ (48)
$\beta_{\ell}^{j}$ =$=\begin{cases}1,&\ell=J_{1}+j\text{,}\\ 0,&\text{otherwise}\text{.}\end{cases}$ (49)
@@ III-A Simulation Results
Figure 1: Source (left) and ground-truth (right) images (original scale). The $401\times 401$ pixel source is a CC0-licensed image, Airy_disk_D65 [40], after conversion to grayscale and division by 255. The ground-truth image, shown after pixel values are squared, is the entrywise square root of the source image, after masking (occulting) symmetric inner and outer regions, defined by two circles (left overlays), and after adding spots to simulate twin exoplanets.
@@
The ground-truth image does not exhibit a phase aberration. To synthesize an aberrated image, a random phase offset is introduced entrywise in the Fourier domain, often called the pupil plane in the context of coronagraphy. Because pixel values in the image plane, the inverse Fourier domain, must be real, this phase aberration is constrained to obey the required symmetry. Notwithstanding a few entries, such as the DC value, each entry of the random offset, an image-like matrix, belongs to a uniform distribution of $-\pi$ to $+\pi$ radians.
@@ IV-A Tensor Objects
For a tensor object, the degree method returns the number of indices, i.e., the numel of the indices property, and the ndims methods returns the number of dimensions, i.e., the ndims of the entries property. Because first and second dimensions are MDA rows and columns, the degree is at least two less than the ndims of a tensor object.
@@ IV-B Index Objects
%Test2 assert(isa(k,’index’))
k = index(2); %Test5

### v4

@@ I Introduction
“Tensors” easily associate with learning-based approaches to imaging. After all, TensorFlow from Google Brain [Abadi2016] and Tensor Comprehensions from Facebook AI [Vasilache2018] are machine learning systems connected with a well known deep learning revolution in object detection and classification. Nonetheless, imaging researchers like Mai et al. [Mai2022] highlight the importance of model-based tensor approaches, emphasizing explanatory value. Mai et al. also argue that a model-based approach helps to reduce the training data requirements for a learning-based approach, yielding what is in effect a hybrid approach.
@@
Imaging researchers are finding advantages of model-based tensor approaches in a variety of applications. These include: multi-exposure, multi-focus, and hyper/multi-spectral image fusion to increase dynamic range, depth-of-field, and spatial resolution [Mai2022, Wang2023, Prevost2022]; inpainting, or tensor completion, of missing or corrupt pixels in images and videos, even rows or columns of the same [Liu2021, Bengua2017]; deblurring, or image restoration, posed as a constrained inverse problem often with added noise [Newman2020, Lefkimmiatis2015, Rezghi2017]; compressive sensing, a mixture of computational imaging and image compression, for light field imaging (LFI), synthetic aperature radar, and hyperspectral imaging modalities [Marquez2020, Qiu2020, Li2020]; other constrained image enhancements, like ones that enforce 3D spherical invariances [Skibbe2017]; and image quality assessments, like ones tailored for LFI [Zhou2020].
@@
Model-based imaging research involving tensors [Mai2022][Zhou2020] exhibits patterns. Papers typically have a section or subsection, on fundamentals, that addresses unary, binary, and $N$-ary operations expressible with tensors, like multi-way or multi-index contractions, Kronecker and Khatri-Rao products, and Tucker, canonical polyadic (CP), and tensor trains (TT) constructs. More fundamental than operations, authors discuss an $n$-mode notation, especially for decomposition constructs, an index notation, also called Einstein notation, or a hybrid notation thereof. Sometimes, extended matrix-vector (EMV) operators are employed to represent Hadamard, or entrywise, and Kronecker, or outer, products. Notation can resemble MATLAB or Python+NumPy syntax for multidimensional arrays. Instead of matrix-vector (MV) or EMV algebra, authors prefer tensor algebra to represent and manipulate high-order image relationships, including generalizations of the singular value decomposition (SVD) for rank reduction and feature extraction, akin to principal components analysis, and optimizations where images are the unknown variables in a functional that suffers upon vectorization of variables.
@@
When literature on learning-based [Abadi2016][Vasilache2018] and model-based [Mai2022][Zhou2020] approaches to imaging use the word, “tensor” mainly or solely means a numeric data structure, usually an MDA, together with numeric operations, like those identified with overloadable operators in object-oriented programming languages. This paper uses the word likewise, avoiding requirements for deeper tensor character addressed, for example by Synge and Schild [Synge1978], in books on geometric tensor calculus.
@@
Whereas tensor notations and operations are in some senses implementation agnostic, a brief survey of tensor software is in order. Because its components, like the Cyclops Tensor Framework (CTF) [Solomonik2014, Singh2022], may apply to tensor contractions, completions, and decompositions for imaging, the NorthWest Chemistry (NWChem) software ecosystem is noteworthy [Valiev2010, Kowalski2021]. With the Tensor Contraction Engine (TCE) and the Tensor Algebra for Many-Body Methods (TAMM) [Baumgartner2005, Mutlu2023], this ecosystem targets massively parallel computing and large-scale molecular simulations. Intended more for imaging and for desktop parallel computing, Tensor Learning in Python (TensorLy) offers tensor decompositions to improve efficiency, robustness, and explainability of deep learning networks [Kossaifi2019, Panagakis2021]. For desktop computing with MATLAB, the Partial Least Squares (PLS) and Multivariate Image Analysis (MIA) Toolboxes, cited in biomedical imaging research [Cumpson2015, Bedia2020], offer tensor decompositions and imaging customizations.
@@
As Harrison and Joseph argue [Harrison2016], the codesign of tensor algebra and software, what they call a tensor framework, has benefits. Codesign facilitates programmatic and computational efficiencies for solving a variety of model-based problems. In support, they cite Bader and Kolda’s [Bader2006, Kolda2009] development of $n$-mode+ notation and the MATLAB Tensor Toolbox (MTT), also called the Tensor Toolbox for MATLAB. While Bader and Kolda contributed notation for decomposition constructs along with original software, others leveraged existing notations while contributing decomposition constructs and MATLAB toolboxes. These include structured data fusion (SDF) with Tensorlab, led by Sorber and De Lathauwer [Sorber2015, Vervliet2014], and the TT-Toolbox with the TT construct, by Osledets [Oseledets2011].
@@
This paper proposes the Ricci-notation tensor (RT) framework for model-based approaches to imaging. A successor to Harrison and Joseph’s numeric tensor (NT) framework [Harrison2016], itself a complement to an EMV framework, the RT framework inherits their advantages. The proposed framework comprises RT algebra and RT software, defined as the RTToolbox [Joseph2023], developed with this paper, plus MATLAB. The RT software parses the RT algebra, dispatching calculations to the MATLAB kernel, with no dependencies on other toolboxes.
@@
The NT software had options [Harrison2016]. With LibNT, C/C++ end users expressed the NT algebra with programmatic and, especially, computational efficiency. With NTToolbox, comprising M-file and MEX wrappers for LibNT, MATLAB end users expressed the NT algebra with some programmatic and some computational penalties. The RT software has only the RTToolbox option [Joseph2023], although this paper elaborates on its design in a C/C++-compatible way. Comprising just M-files, no compilation or linking is required. As per MATLAB Central recommendations [MathWorks2022], avoidance of MEX files and binaries means that sharing is easy via File Exchange.
For its computational efficiency, the RT software leverages compiled unary and binary pagewise functions introduced, in 2020 and 2022, to the MATLAB kernel. Of note, the kernel pagewise multiplication [MathWorks2020] resembles the LibNT lattice multiplication [Harrison2016] introduced, in 2016, with the NT software. This paper elaborates on the lattice concept to facilitate RT left- and right-division using kernel functions. It also introduces pagewise concatenation, an interpreted function the RT software exploits to support $N$-ary tensor concatenation. Programmatic efficiency of the RT software is illustrated via software expressions cross-referenced to algebraic ones.
An image enhancement example, developed here, illustrates the RT algebra. Inspired by Sirbu et al.’s model-based research [Sirbu2013, Sirbu2016, Sirbu2017] for coronagraphy [Schulz2023], the example adopts a simple coronagraph model with two planes related by 2D discrete Fourier transforms and a phase aberration. Starting with a Creative Commons public domain (CC0) image [Wikipedia2022], called Airy_disk_D65, the impact of the model on a synthetic ground-truth image having an occulted star with nearby exoplanets is computed. The exoplanets become visible after correction of the aberration, treated as unknown.
@@ II-B Nonscalars and Broadcasting
Unlike the NT algebra, the RT algebra supports outer additions, subtractions, concatenations, and relations. The unifying idea is that any tensor of one degree may be expanded to an equivalent tensor of higher degree via an outer product with a unit tensor of the difference degree. Consider the following equations and relation, which consistently employ a degree-two matrix, $\mathbf{A}$, four degree-one column-vectors, $\mathbf{1}$, $\mathbf{b}$, $\mathbf{c}$, and $\mathbf{1}$, and an entrywise absolute-value operation, $|\cdot|$:
@@
$\alpha_{\ell}^{j}$ $=\begin{cases}1,&\ell=j\text{,}\\ 0,&\text{otherwise}\text{,}\end{cases}$ (42)
$\beta_{\ell}^{j}$ $=\begin{cases}1,&\ell=J_{1}+j\text{,}\\ 0,&\text{otherwise}\text{.}\end{cases}$ (43)
@@ III-A Simulation Results
Figure 1: Source (left) and ground-truth (right) images (original scale). The $401\times 401$ pixel source is a CC0-licensed image, Airy_disk_D65 [Wikipedia2022], after conversion to grayscale and division by 255. The ground-truth image, shown after pixel values are squared, is the entrywise square root of the source image, after masking (occulting) symmetric inner and outer regions, defined by two circles (left overlays), and after adding spots to simulate twin exoplanets.
@@
The ground-truth image does not exhibit a phase aberration. To synthesize an aberrated image, a random phase offset is introduced entrywise in the Fourier domain, often called the pupil plane in the context of coronagraphy. Because pixel values in the image plane, the inverse Fourier domain, must be real, this phase aberration is constrained to obey the required symmetry. Notwithstanding a few entries, such as the DC value, each entry of the random offset, an image-like matrix, belongs to a uniform distribution of $\pi$ to $\pi$ radians.
@@ IV-A Tensor Objects
For a tensor object, the degree method returns the number of indices, i.e., the numel of the indices property, and the ndims method returns the number of dimensions, i.e., the ndims of the entries property. Because first and second dimensions are MDA rows and columns, the degree is at least two less than the ndims of a tensor object.
@@ IV-B Index Objects
%% Test 2 assert(isa(k,’index’))
k = index(2); %% Test 5